A body thrown up with a finite speed is caught back after 4 sec.the speed of body with which it is thrown up is..

Grade:12th pass

1 Answers

Bidisha Baruah
8 years ago
at the highest point,v=0m/s and g=9.8m/s2,t=4/2sec=2sec
now,v=u+gt
0=u-9.8x2
so,u=19.6m/s